Give the steps of solving the problem:


 If 137+276=435 then how much is 731+672=???????


a.534 b.1403 c.1623 d.1513

Grade:Upto college level

1 Answers

ajinkya
9 years ago
7×100+3×10+1 ? ? ? + 6×100+7×10+2 =1×1000+4×100+0×10+3 =1403.b